OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 − 16 = 0 and x2 = −2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

(x-4)(x+4) = 0 x = 4 or x = -4 x2 + 2x -24 = 0 (x+6)(x-4) = 0 x = -6 or x = 4 The graphing of the solutions shows that they have 1 solution in common: x=4

OpenStudy (anonymous):

\[x^{2}-16=0,x ^{2}=16,x=\pm4.\] \[x^{2}=-2x+24,x ^{2}+2x-24=0,x ^{2}+6x-4x-24=0,\] or x(x+6)-4(x+6)=0 or (x+6)(x-4)=0 Either x+6=0,gives x=-6 Or x-4=0, gives x=4 Solution is x=-6, x=4

OpenStudy (anonymous):

l..L.et f(x)=x^2-16 Again Let f(x)=x^2+2x-24=x^2+2x+1-25 f(x)=(x+1)^2-25 f(x)+25=X^2,which is a parabola with vertex at(-1,-25) Two graphs are parabolas They are alike.